wiki:UC20180604

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 6/04/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Mary Beth Kaiser (JHU), John Bally (CU), Joanne Hughes Clark (Seattle U), Kal Kadlec (UW), Anne Verbiscer (UVa), Ben Williams (UW), Moire Prescott (NMSU), Russet McMillan (APO), Bill Ketzeback (APO), Misty Bentz (GSU)


User feedback and comments from institutional representatives:

  • Washington - nothing to report
  • NMSU - question from a user about ARCTIC - the dark current is high, time variable and displays structure across the chip - are there tips for how to account for this when reducing the data? Russet said that the DC current should not be variable other than scaling w/exposure time. We have to remind users to not apply them to very short exposures, instead people can scale it by hand with individual exposure times. Joanne mentioned that she did take a series of darks last night during cloudy weather; she will analyze it this week and will report how linear the dark current looks. Russet also suggested that Moire have the user contact Russet to describe the issues in more detail.
  • JHU - nothing to report
  • Colorado - nothing to report
  • UVa - nothing to report
  • Georgia State - nothing to report
  • Oklahoma - no report
  • Wyoming - no report
  • Seattle - Working on testing the dark current (see above), otherwise nothing to report.
  • NAPG - no report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 5/1/18 through 5/30/18

0) Overview

Reasonably good weather this period. DIS Blue Camera vacuum was serviced again this past week to address ongoing scattered light issue. We have scheduled all the M2 and M3 mirror work for the summer and a PO was placed with the vendors. We are also busy with job interviews to replace Fritz, who is retiring this summer, and Beth, who will be leaving in the fall.

1) Telescope

Telescope was operational with a couple of issues. We were seeing some oscillations on the telescope and adjusted some of the servo tuning parameters. We had a failure on the secondary that required an extensive work-around that night but thanks to many here at APO and Conor at UW the impact to observing was kept to a minimum.

2) Instruments

DIS: Red camera scattered light is holding well, but blue deteriorated enough to cause us to service it again. Repaired !DCam was reinstalled and is working beautifully. Will send out spare camera for repair soon.

TripleSpec has been operational.

Agile has been operational this period. Rotator issues are still with us and we are looking into whether it has gotten worse and if any improvement can be made.

Echelle is operational. A replacement slit viewer camera order from Fingerlakes has been ordered but not yet received.

NIC-FPS is operational.

ARCTIC is operational. We will be working on improving filter wheel reliability over summer shutdown. Users should verify that their desired filter is in place.

3) ARCSAT

M1 and M2 collimation completed. We have some initial sky data that looks good. We have implemented some secondary software fixes by DFM, however, there is more that is needed and so far we have not seen any real improvement in focus repeatability. The telescope is operational. We had 1 readout issue with SurveyCam that we are monitoring. There are still bugs in the commercial software that will likely always be with us.


Additional telescope and instrument discussion:

Nothing additional to report beyond what was sent out by Mark (see above). Bill and Nancy noted the heroic efforts on the part of the APO staff and Conor at UW for dealing with the secondary issue quickly and efficiently.


Discussion of summer shutdown:

Summer shutdown is scheduled to start for the 3.5m on July 5. On that day the engineering staff will disassemble the telescope. The secondary and tertiary mirrors and the SDSS secondary will leave the site on the following Mon or Tues and will be taken to Tucson, where Mark will drop off the 3.5m tertiary and will then take the other mirrors to the coater in California. The servicing of the mirrors will proceed from there. Other maintenance issues will keep the site staff busy during the remainder of the shutdown. It is unusually long this year (see wiki:UC20180507 minutes from May 2018 telecon for more details) but if things go smoothly at the vendors and transportation of the mirrors goes well then we can possibly end the shutdown early. We are planning a number of telescope and instrument servicings during the shutdown but nothing unusual that we haven’t done multiple times before. John Bally asked if there is any worry that the mirrors won’t get back in time? Bill replied that this is a small but non-zero possibility, and that we would have to extend shutdown if that were to happen. Russet reminded the group that because we are recoating the mirrors and having the tertiary refigured it will be vital to get engineering time on sky, with good weather, before turning the telescope back over to science operations.


Q2 3.5-m scheduling:

Russet is issuing a new Q2 schedule today. All of the unassigned time is now assigned, so there is no remaining available time in Q2.


Q3 3.5-m scheduling:

There are 4 nights in July prior to the shutdown and some of that time was already promised to a Q2 overflow program. For the 4 remaining half nights on July 1-4 there were 8 requests; Russet moved one to June and others are sharing the time. In Q3 there is a higher proportion than normal of partial night programs. Russet arranged some sharing, and made note of some programs that requested specific dates if the shutdown should end early. There are currently 7 half nights available (unassigned) in Q3; A halves are oversubscribed and this oversubscription may extend into Q4 as proposers are chasing the Kepler field.


Q2-3 ARCSAT scheduling:

Ben sent out call for the June+Q3 ARCSAT schedule. Please spread the word that ARCSAT is now operational and performing better than ever!


NA2 Baffle Project

John Wilson provided an email report, which Nancy relayed to the Users Committee. No progress has been made on this project in the past month, and John will be traveling for most of June. Work on this project will resume in July.


KOSMOS update

Users are starting to ask when they can expect to see KOSMOS on the 3.5m telescope. This is still fairly uncertain given that we have not yet received the instrument from NOAO, but if it comes in the next month or so, a very optimistic estimate was by end of the calendar year. The UW team knows what they will need to modify in order to adapt KOSMOS to the 3.5m. The custom shipping crates have been ordered by UW and should be sent to NOAO next week.


TEWG Update

The Training and Education Working Group (TEWG) has had one telecon so far. The survey for current and former students who have been trained on the 3.5m is coming soon; Nancy asked that everyone help by spreading the word and asking students in your departments to complete the survey.


ACTION ITEMS:

Action items from previous meetings:

  • Observers should report to the obs-specs if they notice anything unusual about the telescope tracking accuracy that may be caused by a tracking oscillation in azimuth direction. The effect will likely not be noticeable with DIS but will be detectable with ARCTIC, so not all observers may notice anything. But if you see something, say something. CLOSED: Russet has been watching this closely and she cannot see effects on any integrated measurements on any instruments except APOLLO. Improving oscillations on telemetry does lead to better image quality, but the performance payoff not readily detectable.
  • Institutional schedulers should retain a list of programs that could be inserted into the schedule in late August/early September if the summer shutdown ends up being shorter than anticipated. CLOSED: We received several of these requests from NMSU and UW and 1-2 other institutions with smaller allocations. We would be willing to entertain more from other institutions, especially those that didn’t get a lot of time in Q3.

New action items from this meeting:

None.


Next meeting:

The next meeting was nominally scheduled for July 2. However, we will have no new information about the summer shutdown at that time since we will be about to start shutdown. Therefore the July 2 telecon is canceled. We will resume our telecon schedule the following month: August 6, at 11:00 MDT. At that time we also will examine the telecon schedule and make adjustments if necessary to accommodate fall teaching schedules.

Note: See TracWiki for help on using the wiki.